OpenStudy (anonymous):

3y = 4x If x = 5 in the equation above, y is how much greater than x?

OpenStudy (anonymous):

plug in 5 for x

OpenStudy (anonymous):

y=4x/3 y=4(5)/3=20/3 (20/3)/5=4/3, so y is 4/3 times bigger than x

OpenStudy (anonymous):

@ImAEtrustme let them try before you go giving answers

OpenStudy (anonymous):

@lexiamee obviously they've tried if they're resorting to the internet for answers. If you just type a question like this, I assume you just want an answer. Also, If it's straightforward like this, I assume they can follow my logic. If not, just ask me and I'll clarify.